A block of ice with mass 2.00 kg slides 0.63 m down an inclined plane that slopes downward at an angle of 37.8∘ below the horizontal.

A) If the block of ice starts from rest, what is its final speed? You can ignore friction.
Answer #1

acceleration of block

a = g sin x = 9.8 sin 37.8

a = 6 m/s^2

using 3rd equation of motion

v^2 = 2 a d

v^2 = 2* 6* 0.63

v = 2.75 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck